A football is kicked with an initial velocity of 20 m/s at an angle of 50-degrees with the horizontal. Determine the time of flight, the horizontal distance, and the peak height of the football.
1
Expert's answer
2021-10-15T10:22:25-0400

The time of flight

"T = \frac{2usin θ}{g} \\

= \frac{2 \times 20 \times sin 50°}{9.81} \\

= 3.12 \;s"

The horizontal distance

"R = \frac{u^2sin(2θ)}{g} \\

= \frac{20^2 \times sin(2(50°))}{9.81} \\

= 40.15\;m"

The peak height of the football

"H = \frac{u^2sin^2θ}{2g} \\

= \frac{20^2 \times sin^2(50°)}{2 \times 9.81} \\

= 11.96 \;m"